calculate the fundamental frequency of the wave in an organ pipe of length 4.00cm which is closed at one end. The temperature of the organ pipe is 35 degree celsius.
Answer #1

Solution :

Given :

Length of organ pipe (L) = 4 cm = 0.04 m

Temperature = 35o C

And, Speed of sound at 35o C is 351.81 m/s.

Now, For organ pipe open at one end : f = v / 4L = (351.81 m/s) / {(4)(0.04 m)} = 2198.8 Hz